# PyramidInfer: Pyramid KV Cache Compression for High-throughput LLM Inference

### v1

@@ Abstract
Dongjie Yang1,thanks:   Dongjie Yang and Hai Zhao are with the Department of Computer Science and Engineering, Shanghai Jiao Tong University, and also with Key Laboratory of Shanghai Education Commission for Intelligent Interaction and Cognitive Engineering, Shanghai Jiao Tong University., Xiaodong Han2, Yan Gao2, Yao Hu2, Shilin Zhang3, Hai Zhao1,11footnotemark: 1,thanks:   Corresponding author; This paper was partially supported by Joint Research Project of Yangtze River Delta Science and Technology Innovation Community (No. 2022CSJGG1400). 1 Shanghai Jiao Tong University, 2 Xiaohongshu Inc., 3 South China University of Technology 1{djyang.tony@,zhaohai@cs.}sjtu.edu.cn, 2{shuweng,yadun,xiaohou}@xiaohongshu.com
@@ 3.1.2 Discussion
The inference process differs from training because all the input tokens predict the next tokens. At this time, keys and values store two kinds of information: 1) the information to predict what the token is next to it; 2) the context information for future tokens to leverage. So far, we have verified that PvCs are the crucial keys and values that are useful for inference. On the other hand, we want to verify the non-PvC that may play a more important role in teacher-forcing prediction instead of being the context. As non-PvCs are trivial in PyramidInfer, we discuss it in the Appendix 3.2.2.
@@ 3.2.2 Discussion
The Association between ICR and RAC
Further Verification of ICR about the Role of Non-PvCs
Figure 7: The composition of the keys and values of one layer.
  1. 1.
  2. 2.
  3. 3.
@@ 5.1 Basic Evaluation
In Table 10, we fix the input length and the batch size. For LLaMA 2-13B, PyramidInfer showcases 2.24x throughput than full cache using Accelerate with 54.6% less GPU memory in the KV cache. For LLaMA 2-70B, PyramidInfer can still generate in the prefill phase compared to other me. Existing KV cache compression methods like H2O can not even process the prompt and strike the OOM before the start of compression.

### v2

@@ Abstract
Dongjie Yang1,thanks:   Dongjie Yang and Hai Zhao are with the Department of Computer Science and Engineering, Shanghai Jiao Tong University; Key Laboratory of Shanghai Education Commission for Intelligent Interaction and Cognitive Engineering, Shanghai Jiao Tong University; Shanghai Key Laboratory of Trusted Data Circulation and Governance in Web3., Xiaodong Han2, Yan Gao2, Yao Hu2, Shilin Zhang3, Hai Zhao1,11footnotemark: 1,thanks:   Corresponding author; This paper was partially supported by Joint Research Project of Yangtze River Delta Science and Technology Innovation Community (No. 2022CSJGG1400). 1 Shanghai Jiao Tong University, 2 Xiaohongshu Inc., 3 South China University of Technology 1{djyang.tony@,zhaohai@cs.}sjtu.edu.cn, 2{shuweng,yadun,xiahou}@xiaohongshu.com
@@ 3.1.2 Discussion
The inference process differs from training because all the input tokens predict the next tokens. At this time, keys and values store two kinds of information: 1) the information to predict what the token is next to it; 2) the context information for future tokens to leverage. So far, we have verified that PvCs are the crucial keys and values that are useful for inference. On the other hand, we want to verify the non-PvC that may play a more important role in teacher-forcing prediction instead of being the context. As non-PvCs are trivial in PyramidInfer, we discuss it in the Appendix B.
@@ 5.1 Basic Evaluation
In Table 8, we fix the input length and the batch size. For LLaMA 2-13B, PyramidInfer showcases 2.24x throughput than full cache using Accelerate with 54.6% less GPU memory in the KV cache. For LLaMA 2-70B, PyramidInfer can still generate in the prefill phase compared to other methods. Existing KV cache compression methods like H2O can not even process the prompt and strike the OOM before the start of compression.
@@ A.2 Position Encoding
Appendix B Extended Discussions
The Association between ICR and RAC
Further Verification of ICR about the Role of Non-PvCs
Figure 10: The composition of the keys and values of one layer.
  1. 1.
  2. 2.
  3. 3.